# On the Lebesgue Constant of Extended-Domain Spectral Methods for Elliptic PDEs

### v2

@@
Abstract
The extended-domain method is a simple and effective strategy for applying spectral methods to complex geometries. Its stability, however, is complicated by the use of a Fourier extension basis, which is known to be exponentially ill-conditioned. Numerical experiments reveal a discrepancy: the method appears stable for the non-self-adjoint convection-diffusion equation but unstable for the self-adjoint Poisson equation. This paper provides a rigorous analysis that resolves this issue. We first prove that the method is asymptotically unstable for both operators; the Lebesgue constant is shown to grow exponentially as a direct consequence of the frame redundancy of the basis. We then explain the observed discrepancy as a pre-asymptotic phenomenon. We prove a structural dichotomy in the discrete Green’s functions (the inverses of the physical-space operators): the inverse of the convection-diffusion operator is numerically quasi-sparse, with entries that decay exponentially off-diagonal, in stark contrast to the numerically dense inverse of the Poisson operator. This framework provides a rigorous theoretical foundation that reconciles the method’s practical utility in certain regimes with its ultimate asymptotic instability.
keywords:
@@ Abstract
journal: Applied and Computational Harmonic Analysis
1 Introduction
Spectral methods are widely used in scientific computing due to their potential for exponential convergence rates, although their direct application is typically restricted to problems on simple tensor-product domains [1, 2]. A principal challenge is the extension of these methods to handle complex geometries. A variety of “fictitious domain” or “domain embedding” methods have been developed to address this issue, including methods based on interface tracking [3], smooth solution extension [4, 5], penalization [6], and the immersed boundary method [7].
One of the simplest strategies is the extended-domain method, wherein a global basis is defined on a larger, regular domain that contains the physical one [8, 9, 10, 11]. The stability of this approach, however, has remained a topic of concern. The underlying basis, a Fourier series on an extended domain, is known to form a frame, not an orthonormal basis, on the physical sub-domain, leading to a notoriously ill-conditioned function approximation problem [12, 13]. When this basis is used to discretize a differential operator, the ill-conditioning is amplified [14, 15]. Curiously, numerical evidence suggests a far better-conditioned system for the non-self-adjoint convection-diffusion equation than for the self-adjoint Poisson equation. This apparent contradiction between the known instability of the basis and the observed stability for certain operators motivates the present investigation.
This paper provides a rigorous stability analysis that reconciles the asymptotic theory with pre-asymptotic numerical observations. The primary contributions are as follows:
  • 1.
    We rigorously prove that the method is asymptotically unstable for both the Poisson and convection-diffusion operators. We demonstrate that the Lebesgue constant grows exponentially, a direct consequence of the severe ill-conditioning of the underlying Fourier extension frame.
  • 2.
@@ 1 Introduction
    We explain the observed pre-asymptotic stability of the convection-diffusion operator by proving a fundamental structural dichotomy in the discrete Green’s functions. We rigorously establish that the inverse of the discrete physical-space convection-diffusion operator is numerically quasi-sparse due to exponential off-diagonal decay, in stark contrast to the numerically dense inverse of the Poisson operator.
  • 3.
@@
    We identify the modal Péclet number as the physical mechanism governing the transition from the well-behaved pre-asymptotic regime to the true asymptotic instability, providing a multiscale picture of the method’s performance.
These results provide a rigorous theoretical framework for this practical method, explaining why it can be effective in certain convection-dominated regimes while confirming its ultimate theoretical limitations. The remainder of this paper is organized as follows. Section 2 establishes the spectral collocation framework. Section 3 proves the asymptotic exponential instability. In Section 4, we prove the pre-asymptotic stability properties of the convection-diffusion operator. Section 5 extends these findings to higher dimensions. Section 6 validates our theoretical framework with numerical experiments. Finally, Section 7 offers concluding remarks.
2 The Spectral Collocation Framework on an Extended Domain
@@ 2.1 Basis and Discretization
$w^{\prime\prime}=\lambda_{j}w\quad\text{in }\tilde{I},\quad w(-\delta)=w(L+\delta)=0.$ (1)
@@
$w_{j}(x)=\sqrt{\frac{2}{L+2\delta}}\sin\left(\frac{j\pi(x+\delta)}{L+2\delta}\right),\quad\lambda_{j}=\left(\frac{j\pi}{L+2\delta}\right)^{2},\quad j\in\mathbb{N}.$ (2)
@@
The solution is approximated in the finite-dimensional space $V_{N}=\text{span}\{w_{1},\dots,w_{N}\}$ by an expansion $u_{N}(x)=\sum_{j=1}^{N}c_{j}w_{j}(x)$. Given a differential operator $\mathcal{L}$, the functions $\{\phi_{j}(x)=\mathcal{L}[w_{j}](x)\}_{j=1}^{N}$ form the interpolation basis. For a set of $N$ distinct collocation points $\{x_{k}\}_{k=1}^{N}\subset I$, the coefficients $\mathbf{c}=(c_{1},\dots,c_{N})^{T}$ are found by enforcing the PDE at these points:
$\mathcal{L}[u_{N}](x_{k})=\sum_{j=1}^{N}c_{j}\phi_{j}(x_{k})=f(x_{k}),\quad k=1,\dots,N.$ (3)
This defines an $N\times N$ linear system $\mathbf{A}\mathbf{c}=\mathbf{f}$ where the collocation matrix entries are $\mathbf{A}_{kj}=\phi_{j}(x_{k})$.
@@
2.2 Stability and the Lebesgue Constant
The stability of the collocation scheme is determined by the conditioning of the matrix $\mathbf{A}$. A robust and basis-independent measure of this stability is the Lebesgue constant of the underlying interpolation operator.
@@ 2.2 Stability and the Lebesgue Constant
$\Lambda_{N}=\sup_{g\in C(I),g\neq 0}\frac{\left\lVert\mathcal{I}_{N}g\right\rVert_{L^{\infty}(I)}}{\left\lVert g\right\rVert_{L^{\infty}(I)}}.$ (4)
@@
$\Lambda_{N}=\max_{x\in I}\sum_{k=1}^{N}\left\lvert l_{k}(x)\right\rvert$. (5)
A Lebesgue constant $\Lambda_{N}$ that grows polynomially with $N$ ensures the stability and, ultimately, the convergence of the method. The central analysis of this paper is the rigorous bounding of this quantity for the Poisson and convection-diffusion operators.
3 Asymptotic Instability for the Poisson Equation
In this section, we establish the asymptotic stability baseline for the extended-domain method. We rigorously prove that, when applied to the self-adjoint Poisson equation, the method is fundamentally unstable, with a Lebesgue constant that grows exponentially with the number of modes $N$. This instability is a direct and unavoidable consequence of the exponential ill-conditioning inherent in the underlying Fourier extension approximation.
The core of the problem lies in the properties of the basis functions $\{w_{j}(x)\}_{j=1}^{N}$ defined on the extended domain $\tilde{I}=(-\delta,L+\delta)$ when they are restricted to and collocated within the physical domain $I=(0,L)$. On $L^{2}(I)$, this set of functions does not form an orthogonal or even a stable Riesz basis. Instead, it constitutes a frame that is highly redundant. This redundancy leads to near-linear dependencies among the high-frequency basis functions when sampled on $I$, which is the source of the exponential ill-conditioning.
Theorem 3.1 (Exponential Growth of the Lebesgue Constant for the Poisson Operator).
Let the extended-domain spectral collocation method be applied to the Poisson equation $u^{\prime\prime}=f$. The associated Lebesgue constant, $\Lambda_{N}$, grows exponentially with $N$. Specifically, there exist constants $C,\alpha>0$, independent of $N$, such that for sufficiently large $N$:
$\Lambda_{N}\geq Ce^{\alpha N}.$
@@ 3 Asymptotic Instability for the Poisson Equation
The proof proceeds by constructing a specific coefficient vector $\mathbf{c}$ for which the norm of the resulting function, $u_{N}(x)=\sum_{j=1}^{N}c_{j}w_{j}(x)$, is exponentially small at the collocation points inside the physical domain $I=(0,L)$, while the norm of $\mathbf{c}$ itself is $O(1)$. This directly demonstrates the exponential ill-conditioning of the collocation matrix $\mathbf{W}$.
[…]
2. Bounding the Function on the Physical Domain. The function $u_{N}(x)$ is analytic and approximates $g(x)$, which is identically zero for $x\in I$. For such approximations, it is a classical result in approximation theory that the function $u_{N}(x)$ must be exponentially small throughout the domain $I$, far from the support of $g(x)$ (see, e.g., [16, Sec. 2.1.4]):
$\max_{x\in I}|u_{N}(x)|\leq C_{1}e^{-\alpha_{1}N}$
for some constants $C_{1},\alpha_{1}>0$ that depend on the geometry ($L,\delta$) but not on $N$.
3. Bounding the Smallest Singular Value of $\mathbf{W}$. Let $\mathbf{u}_{N}$ be the vector of values of $u_{N}(x)$ at the collocation points $\{x_{k}\}_{k=1}^{N}\subset I$. From the previous step, $\left\lVert\mathbf{u}_{N}\right\rVert_{\infty}\leq C_{1}e^{-\alpha_{1}N}$. By definition, $\mathbf{u}_{N}=\mathbf{W}\mathbf{c}$. The smallest singular value of $\mathbf{W}$ is $\sigma_{\min}(\mathbf{W})=\min_{\mathbf{v}\neq 0}\frac{\left\lVert\mathbf{W}\mathbf{v}\right\rVert_{2}}{\left\lVert\mathbf{v}\right\rVert_{2}}$. Using our constructed vector $\mathbf{c}$:
$\sigma_{\min}(\mathbf{W})\leq\frac{\left\lVert\mathbf{W}\mathbf{c}\right\rVert_{2}}{\left\lVert\mathbf{c}\right\rVert_{2}}=\frac{\left\lVert\mathbf{u}_{N}\right\rVert_{2}}{\left\lVert\mathbf{c}\right\rVert_{2}}$
Using the norm equivalence $\left\lVert\mathbf{v}\right\rVert_{2}\leq\sqrt{N}\left\lVert\mathbf{v}\right\rVert_{\infty}$ and the fact that $\left\lVert\mathbf{c}\right\rVert_{2}$ is bounded below by a constant, we get:
$\sigma_{\min}(\mathbf{W})\leq\frac{\sqrt{N}\left\lVert\mathbf{u}_{N}\right\rVert_{\infty}}{\left\lVert\mathbf{c}\right\rVert_{2}}\leq C_{2}\sqrt{N}e^{-\alpha_{1}N}$
The exponential term dominates the polynomial term, so $\sigma_{\min}(\mathbf{W})$ decays exponentially with $N$.
4. Bounding the Condition Number and Lebesgue Constant. The collocation matrix for the Poisson operator is $\mathbf{A}=\mathbf{W}\mathbf{\Lambda}$. The norm of its inverse is $\left\lVert\mathbf{A}^{-1}\right\rVert_{2}=\left\lVert\mathbf{\Lambda}^{-1}\mathbf{W}^{-1}\right\rVert_{2}$. Using the properties of singular values, we have:
$\left\lVert\mathbf{A}^{-1}\right\rVert_{2}\geq\sigma_{\min}(\mathbf{\Lambda}^{-1})\left\lVert\mathbf{W}^{-1}\right\rVert_{2}=\frac{1}{\lambda_{N}}\frac{1}{\sigma_{\min}(\mathbf{W})}\geq\frac{1}{C_{3}N^{2}}\frac{e^{\alpha_{1}N}}{C_{2}\sqrt{N}}\geq C_{4}N^{-2.5}e^{\alpha_{1}N}.$
The exponential growth of the Lebesgue constant is a direct consequence, as established by the following lemma. ∎
Lemma 3.2 (A Lower Bound for the Lebesgue Constant).
Let $\mathbf{A}$ be the $N\times N$ spectral collocation matrix. The Lebesgue constant $\Lambda_{N}$ is bounded below by the spectral norm of the inverse matrix as follows:
$\Lambda_{N}\geq\frac{\sigma_{\min}(\mathcal{S}_{N})}{\sqrt{L}}\left\lVert\mathbf{A}^{-1}\right\rVert_{2}$
where $L$ is the length of the domain $I$, and $\sigma_{\min}(\mathcal{S}_{N})$ is the smallest singular value of the continuous synthesis operator $\mathcal{S}_{N}:\mathbb{C}^{N}\to L^{2}(I)$ that maps a coefficient vector $\mathbf{c}$ to the function $u_{N}(x)=\sum_{j=1}^{N}c_{j}\phi_{j}(x)$.
Proof.
The proof is constructive. We build a specific function $g(x)$ whose nodal values $\mathbf{g}$ require a coefficient vector $\mathbf{c}$ of maximum possible norm, and then lower-bound the ratio of the output to input function norms.
1. Constructing the Worst-Case Nodal Data. Let $\mathbf{g}^{*}$ be a right singular vector corresponding to the largest singular value of $\mathbf{A}^{-1}$, which is $\sigma_{\max}(\mathbf{A}^{-1})=\left\lVert\mathbf{A}^{-1}\right\rVert_{2}$. We normalize it such that $\left\lVert\mathbf{g}^{*}\right\rVert_{2}=1$. The corresponding coefficient vector is $\mathbf{c}^{*}=\mathbf{A}^{-1}\mathbf{g}^{*}$, and its norm is precisely:
$\left\lVert\mathbf{c}^{*}\right\rVert_{2}=\left\lVert\mathbf{A}^{-1}\mathbf{g}^{*}\right\rVert_{2}=\left\lVert\mathbf{A}^{-1}\right\rVert_{2}\left\lVert\mathbf{g}^{*}\right\rVert_{2}=\left\lVert\mathbf{A}^{-1}\right\rVert_{2}.$
2. Defining the Input Function $g(x)$. We define a continuous function $g(x)$ whose nodal values at the collocation points are given by $\mathbf{g}^{*}$. We can construct such a function (e.g., via piecewise linear interpolation) so that its maximum value does not exceed the maximum nodal value. Thus, $\left\lVert g\right\rVert_{L^{\infty}(I)}=\left\lVert\mathbf{g}^{*}\right\rVert_{\ell_{\infty}}$. Using the standard vector norm equivalence, we can bound this:
$\left\lVert g\right\rVert_{L^{\infty}(I)}=\left\lVert\mathbf{g}^{*}\right\rVert_{\ell_{\infty}}\leq\left\lVert\mathbf{g}^{*}\right\rVert_{2}=1.$
3. Finding a Lower Bound on the Output Function $u_{N}(x)$. The resulting spectral interpolant is $u_{N}(x)=\mathcal{I}_{N}g(x)=\sum_{j=1}^{N}c^{*}_{j}\phi_{j}(x)$. We establish a lower bound on its $L^{\infty}(I)$ norm by relating it to the $L^{2}(I)$ norm. The synthesis operator $\mathcal{S}_{N}$ maps $\mathbf{c}^{*}$ to $u_{N}(x)$, so its $L^{2}$ norm is bounded by the smallest singular value of $\mathcal{S}_{N}$:
$\left\lVert u_{N}\right\rVert_{L^{2}(I)}=\left\lVert\mathcal{S}_{N}\mathbf{c}^{*}\right\rVert_{2}\geq\sigma_{\min}(\mathcal{S}_{N})\left\lVert\mathbf{c}^{*}\right\rVert_{2}.$
Using the inverse Cauchy-Schwarz inequality for functions, $\left\lVert u_{N}\right\rVert_{L^{\infty}(I)}\geq\frac{1}{\sqrt{L}}\left\lVert u_{N}\right\rVert_{L^{2}(I)}$, we get:
$\left\lVert u_{N}\right\rVert_{L^{\infty}(I)}\geq\frac{1}{\sqrt{L}}\sigma_{\min}(\mathcal{S}_{N})\left\lVert\mathbf{c}^{*}\right\rVert_{2}=\frac{\sigma_{\min}(\mathcal{S}_{N})}{\sqrt{L}}\left\lVert\mathbf{A}^{-1}\right\rVert_{2}.$
4. Assembling the Final Inequality. The Lebesgue constant is defined as the supremum of the ratio $\left\lVert u_{N}\right\rVert_{L^{\infty}(I)}/\left\lVert g\right\rVert_{L^{\infty}(I)}$. For our constructed function $g(x)$, this ratio is bounded below:
$\Lambda_{N}\geq\frac{\left\lVert u_{N}\right\rVert_{L^{\infty}(I)}}{\left\lVert g\right\rVert_{L^{\infty}(I)}}\geq\frac{\frac{\sigma_{\min}(\mathcal{S}_{N})}{\sqrt{L}}\left\lVert\mathbf{A}^{-1}\right\rVert_{2}}{1}=\frac{\sigma_{\min}(\mathcal{S}_{N})}{\sqrt{L}}\left\lVert\mathbf{A}^{-1}\right\rVert_{2}.$
This completes the proof. ∎
4 Pre-Asymptotic Stability of the Convection-Diffusion Operator

### v3

# On the Pre-Asymptotic Stability and Inverse Structure of Extended-Domain Spectral Methods

@@
Abstract.
The extended-domain method is a strategy for applying spectral methods to complex geometries. Its stability is complicated by the ill-conditioning of the Fourier extension frame. This paper provides a rigorous analysis of the method’s pre-asymptotic behavior. We confirm that the spectral collocation system is asymptotically ill-conditioned for both the Poisson and convection-diffusion operators, driven by the redundancy of the underlying frame. However, we prove a fundamental structural dichotomy in their discrete Green’s functions. We show that the inverse of the convection-diffusion operator is numerically quasi-sparse, exhibiting exponential off-diagonal decay, in stark contrast to the numerically dense inverse of the Poisson operator. This intrinsic sparsity explains why the convection-diffusion operator is significantly more robust to the underlying frame instability in practical computations.
Key words and phrases:
@@ Abstract.
1. Introduction
Spectral methods are a cornerstone of scientific computing, prized for their ability to achieve exponential convergence rates for smooth functions [boyd2001chebyshev, canuto2007spectral]. However, their classical formulation relies heavily on tensor-product grids, making their application to complex, non-rectangular geometries challenging. To bridge this gap, various “fictitious domain” strategies have been proposed, including interface tracking [leveque1994immersed], penalization methods [angot1999penalization], and smooth extension techniques [bruno2010high, stein2016immersed].
Among these strategies, the extended-domain method (also known as Fourier extension) is conceptually attractive due to its simplicity. The method embeds the complex physical domain within a larger, regular computational box and approximates the solution using a Fourier series defined on this extended domain [lui2009spectral, sabetghadam2009fourier, fang2011towards].
Despite its utility, the numerical stability of the extended-domain method is a subject of nuance and ongoing research. The fundamental difficulty lies in the approximation properties of the basis. When a Fourier series on an extended domain is restricted to a subset (the physical domain), the basis functions form a frame rather than an orthonormal basis. This redundancy leads to an interpolation problem that is notoriously ill-conditioned. Indeed, the seminal work of Platte et al. [platte2011impossibility] established that for equispaced data, no method can simultaneously achieve geometric convergence and stability; the condition number must grow exponentially if the convergence is geometric. To mitigate this, standard practice often employs oversampling (where the number of collocation points $M$ exceeds the number of modes $N$) or regularization to stabilize the frame [huybrechs2010fourier, adcock2014numerical].
However, a curious discrepancy persists in the pre-asymptotic regime, particularly when the system is square ($M=N$) or only mildly oversampled. Numerical evidence suggests that the method is significantly more robust for the non-self-adjoint convection-diffusion equation than for the self-adjoint Poisson equation. While the underlying basis is unstable in both cases, the convection-diffusion operator appears to suppress this instability in practical computations, whereas the Poisson operator does not. This resonates with the observations of Trefethen et al. [Trefethen1999] regarding non-normal operators, where standard eigenvalue analysis often fails to predict transient behavior.
This paper provides a rigorous structural analysis that resolves this apparent paradox. We investigate the spectral collocation method in the ”square” limit ($M=N$) to isolate the interaction between the differential operator and the basis ill-conditioning. While regularization is necessary for practical solvers, analyzing the square limit acts as a magnifying glass, revealing the intrinsic structural differences in how the operators propagate the frame’s inherent instability. Our contributions are as follows:
  • We characterize the method’s baseline instability. Consistent with the theory of Fourier frames [adcock2014numerical, platte2011impossibility], we demonstrate that the condition number of the linear system grows exponentially. This confirms that the ultimate source of instability is the redundancy of the basis, which affects all operators asymptotically.
@@ 1. Introduction
  • We explain the observed pre-asymptotic stability of the convection-diffusion operator by proving a fundamental structural dichotomy in the discrete Green’s functions (the inverses of the physical-space operators). We rigorously establish that the inverse of the discrete convection-diffusion operator is numerically quasi-sparse, exhibiting exponential off-diagonal decay (visualized later in Figure 2). This is in stark contrast to the numerically dense inverse of the Poisson operator.
@@
  • We identify the modal Péclet number as the mechanism governing the transition from the stable pre-asymptotic regime to the asymptotic instability, providing a multiscale explanation for why convection-dominated problems resist the ill-conditioning of the underlying frame for longer.
These results provide a theoretical foundation that reconciles the method’s known asymptotic limitations with its practical effectiveness in convection-dominated regimes.
2. The Spectral Collocation Framework on an Extended Domain
@@ 2.1. Basis and Discretization
$w^{\prime\prime}=\lambda_{j}w\quad\text{in }\tilde{I},\quad w(-\delta)=w(L+\delta)=0.$ (2.1)
@@
$w_{j}(x)=\sqrt{\frac{2}{L+2\delta}}\sin\left(\frac{j\pi(x+\delta)}{L+2\delta}\right),\quad\lambda_{j}=\left(\frac{j\pi}{L+2\delta}\right)^{2},\quad j\in\mathbb{N}.$ (2.2)
Figure 1. Schematic of the Extended-Domain Method in 1D. The complex physical problem on $I=[0,L]$ is embedded into a larger, regular computational domain $\tilde{I}=[-\delta,L+\delta]$. The solution is approximated by Fourier-like basis functions (e.g., $w_{j}(x)$) that are defined on the full domain $\tilde{I}$ and vanish at the extended boundaries $\delta$ and $L+\delta$, but generally do not satisfy boundary conditions at the physical endpoints 0 and $L$.
@@
The solution is approximated in the finite-dimensional space $V_{N}=\text{span}\{w_{1},\dots,w_{N}\}$ by an expansion $u_{N}(x)=\sum_{j=1}^{N}c_{j}w_{j}(x)$. Given a differential operator $\mathcal{L}$, the functions $\{\phi_{j}(x)=\mathcal{L}[w_{j}](x)\}_{j=1}^{N}$ form the interpolation basis. For a set of $N$ distinct collocation points $\{x_{k}\}_{k=1}^{N}\subset I$, the coefficients $\bm{c}=(c_{1},\dots,c_{N})^{T}$ are found by enforcing the PDE at these points:
$\mathcal{L}[u_{N}](x_{k})=\sum_{j=1}^{N}c_{j}\phi_{j}(x_{k})=f(x_{k}),\quad k=1,\dots,N.$ (2.3)
This defines an $N\times N$ linear system $\bm{A}\bm{c}=\bm{f}$ where the collocation matrix entries are $\bm{A}_{kj}=\phi_{j}(x_{k})$.
@@
2.2. Stability and the Lebesgue Constant
The stability of the collocation scheme is determined by the conditioning of the matrix $\bm{A}$. A robust and basis-independent measure of this stability is the Lebesgue constant of the underlying interpolation operator.
@@ 2.2. Stability and the Lebesgue Constant
$\Lambda_{N}=\sup_{g\in C(I),g\neq 0}\frac{\left\lVert\mathcal{I}_{N}g\right\rVert_{L^{\infty}(I)}}{\left\lVert g\right\rVert_{L^{\infty}(I)}}.$ (2.4)
@@
$\Lambda_{N}=\max_{x\in I}\sum_{k=1}^{N}\left\lvert l_{k}(x)\right\rvert$. (2.5)
Remark 2.4 (Lebesgue Constant as an Instability Detector).
The Lebesgue constant relates approximation stability to algebraic stability via a lower bound. Specifically, for the synthesis matrix $\bm{W}$, it holds that $\kappa(\bm{W})\geq\Lambda_{N}/\sqrt{N}$ [adcock2014numerical]. Recalling the factorization $\bm{A}_{P}=\bm{W}\mathbf{\Lambda}$ from Section 3, an exponential growth in $\Lambda_{N}$ necessitates an exponential growth in the system condition number $\kappa(\bm{A}_{P})$. Thus, while a small $\Lambda_{N}$ does not guarantee a well-conditioned matrix, a large $\Lambda_{N}$ serves as a sufficient condition to prove algebraic instability.
A Lebesgue constant $\Lambda_{N}$ that grows polynomially with $N$ is a necessary condition for the convergence of the method, ensuring that the interpolation error does not diverge for smooth functions. In the context of this paper, we use $\Lambda_{N}$ to empirically confirm the baseline failure of the Poisson operator (Section 6.1). The algebraic stability of the Convection-Diffusion operator, which occurs despite the potential for frame ill-conditioning, is driven by the quasi-sparse inverse structure proven in Section 4.
3. Baseline Instability of the Poisson Operator
In this section, we establish the baseline stability properties of the extended-domain method when applied to the standard Poisson equation. We focus on the case of a square system ($M=N$), where the number of collocation points equals the number of basis modes. While oversampling is known to stabilize such frames [adcock2014numerical], the square case serves as a critical baseline to understand the inherent interaction between the basis redundancy and the differential operator.
Consistent with the impossibility results for equispaced data [platte2011impossibility], we show that the linear system for the Poisson operator is exponentially ill-conditioned. This ill-conditioning is intrinsic to the Fourier extension frame and provides the necessary context for the structural dichotomy observed in the convection-diffusion operator in Section 4.
3.1. Structure of the Collocation Matrix
Recall that the collocation matrix $\bm{A}$ maps the coefficients $\bm{c}$ of the expansion $u_{N}=\sum c_{j}w_{j}$ to the values of the operator $\mathcal{L}[u_{N}]$ at the collocation points. For the Poisson operator $\mathcal{L}=-d^{2}/dx^{2}$, the basis functions are eigenfunctions, $\mathcal{L}[w_{j}]=\lambda_{j}w_{j}$. This allows us to factor the collocation matrix as:
$\bm{A}_{P}=\bm{W}\mathbf{\Lambda},$ (3.1)
where:
  • $\mathbf{\Lambda}=\text{diag}(\lambda_{1},\dots,\lambda_{N})$ is the diagonal matrix of eigenvalues, with $\lambda_{j}\propto j^{2}$.
  • $\bm{W}$ is the synthesis matrix (or frame matrix) with entries $W_{kj}=w_{j}(x_{k})$. It maps the coefficients to the function values on the physical grid.
The stability of the numerical solution $\bm{c}=\bm{A}_{P}^{-1}\bm{f}$ is governed by the condition number $\kappa(\bm{A}_{P})=\left\lVert\bm{A}_{P}\right\rVert_{2}\left\lVert\bm{A}_{P}^{-1}\right\rVert_{2}$.
3.2. Exponential Growth of the Condition Number
The instability of the method arises from the properties of the synthesis matrix $\bm{W}$. The set of basis functions $\{w_{j}\}_{j=1}^{N}$ restricted to the physical domain $I\subset\tilde{I}$ forms a truncated Fourier extension frame. A defining characteristic of such frames is the rapid decay of their singular values.
Theorem 3.1 (Exponential Ill-Conditioning of the Poisson System).
Let $\bm{A}_{P}$ be the $N\times N$ spectral collocation matrix for the Poisson equation on the physical domain $I\subset\tilde{I}$. The condition number $\kappa(\bm{A}_{P})$ grows exponentially with $N$. Specifically, there exist constants $C,\alpha>0$ such that:
$\kappa(\bm{A}_{P})\geq Ce^{\alpha N}.$
@@ 3.2. Exponential Growth of the Condition Number
The condition number is bounded below by the norm of the inverse: $\kappa(\bm{A}_{P})\geq\left\lVert\bm{A}_{P}^{-1}\right\rVert_{2}$. Using the factorization in (3.1), the inverse is $\bm{A}_{P}^{-1}=\mathbf{\Lambda}^{-1}\bm{W}^{-1}$. We can bound the spectral norm of the product from below:
$\left\lVert\bm{A}_{P}^{-1}\right\rVert_{2}=\left\lVert\mathbf{\Lambda}^{-1}\bm{W}^{-1}\right\rVert_{2}\geq\sigma_{\min}(\mathbf{\Lambda}^{-1})\left\lVert\bm{W}^{-1}\right\rVert_{2}.$
We analyze the two terms separately:
  1. (1)
    The Differential Operator (Polynomial Scaling): The matrix $\mathbf{\Lambda}^{-1}$ contains the inverse eigenvalues. Its smallest singular value corresponds to the largest eigenvalue $\lambda_{N}$:
    $\sigma_{\min}(\mathbf{\Lambda}^{-1})=\frac{1}{\lambda_{N}}\propto\frac{1}{N^{2}}.$
  2. (2)
    The Frame Matrix (Exponential Scaling): The term $\left\lVert\bm{W}^{-1}\right\rVert_{2}$ is the reciprocal of the smallest singular value of the synthesis matrix $\bm{W}$, i.e., $\left\lVert\bm{W}^{-1}\right\rVert_{2}=1/\sigma_{\min}(\bm{W})$. It is a well-established result in the theory of Fourier extensions (see [adcock2014numerical, Sec. 4] and [huybrechs2010fourier]) that the singular values of the synthesis operator for a Fourier extension frame decay exponentially to zero. For the discrete matrix $\bm{W}$ in the square case ($M=N$), this implies:
    $\sigma_{\min}(\bm{W})\leq C_{1}e^{-\alpha N},$
    for some $\alpha>0$ determined by the extension ratio. Consequently, the norm of the inverse grows exponentially:
    $\left\lVert\bm{W}^{-1}\right\rVert_{2}\geq\frac{1}{C_{1}}e^{\alpha N}.$
Combining these estimates, we obtain:
$\left\lVert\bm{A}_{P}^{-1}\right\rVert_{2}\geq\frac{1}{C_{2}N^{2}}e^{\alpha N}.$
Since the exponential growth $e^{\alpha N}$ dominates the polynomial decay $N^{-2}$, the condition number grows exponentially. Q.E.D.
Remark 3.2 (The Implication of Square Systems).
The exponential growth of $\kappa(\bm{A}_{P})$ confirms that for $M=N$, the method is numerically unstable; entries in the coefficient vector $\bm{c}$ can become arbitrarily large to represent $O(1)$ functions, leading to catastrophic cancellation errors. While this can be mitigated by oversampling ($M>N$) to control $\sigma_{\min}(\bm{W})$, the analysis of the square system reveals the ”native” behavior of the operator-basis interaction. This baseline instability makes the results of the following section—where the convection-diffusion operator exhibits a completely different structural behavior—all the more striking.
4. Pre-Asymptotic Stability of the Convection-Diffusion Operator